OpenStudy (anonymous):

slope-intercept inequality: if the graph has points 0,4 and -3,0 what is the slope-intercept inequality

OpenStudy (anonymous):

Evaluate x - 2y - 5z when x = -3, y = 1, and z = -2 Answer 15 -5 5 -15

OpenStudy (kira_yamato):

@Coleman slope = 4/-3 = -4/3 Eq'n: y = 4 - 4/3 x

OpenStudy (kira_yamato):

5